OpenStudy (anonymous):

GIVING A MEDAL Find an equation for the nth term of the arithmetic sequence. 9, 11, 13, 15, ...

OpenStudy (anonymous):

nth term=9+(n-1)2 =9+2n-2 =7+2n

OpenStudy (aravindg):

@svaibhavi happy to see you are willing to help fellow users but just make note that you participate the asker in solving the question ,ie guide the asker to the answer instead of giving away an answer ..that is more helpful :)

OpenStudy (aravindg):

refer Openstudy COC http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

gotcha sorry!

OpenStudy (anonymous):

@svaibhavi That is close but my answers dont show that is correct.. can you explain? @AravindG can you help me?